OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. 14th term = -33 15th term = 9

OpenStudy (anonymous):

a+13d=-33 a+14d=9 where a is first term and d is common difference. Now, for nth term use a+(n-1)d. Put values of a and d from the above two equations.

OpenStudy (anonymous):

d = -24 ... ???

OpenStudy (anonymous):

I'm not quite following

OpenStudy (anonymous):

d=9-(-33)=42

OpenStudy (anonymous):

ohh gotcha!

OpenStudy (anonymous):

so d = 42 so the 14th term is -33 -33 = a + (14 -1)42 -33 = a + (13)42 -33 = a + 546 -33 - 546 = a -579 check: 15th term (n = 15) -579+ (15 - 1)42 = ? -579 + (14)42 = ? -579 + (588) = ? =9 so a = -579 d = 42 equation is -579 + 42(n-1)

OpenStudy (anonymous):

@BPDlkeme234 wow! I really appreciate that!

OpenStudy (anonymous):

Elementary! LOL